OpenStudy (anonymous):

Help please :) Use the theorem on inverse functions to prove that f and g are inverse functions of each other, and sketch the graphs of f and g on the same coordinate plane. f(x) = 3x-2 g(x)= (x+2)/3

OpenStudy (anonymous):

so they're not inverses?

OpenStudy (anonymous):

They are inverses. If you plug in g(x) into f(x) you get x. If you get x, then they are inverses. \[f(g(x))= 3[(x+2)/3]-2\]\[f(g(x))= x+2-2\]\[f(g(x))=x\]